OpenStudy (anonymous):

What's 171 / 52 = ? (using an improper fraction as the answer)

hartnn (hartnn):

what is 52*3=?

OpenStudy (anonymous):

156 can you help me solve this equation lol I'm stuck again 28x + 7 + 24x + 18= 180

hartnn (hartnn):

52x=180-18-7 x=52/155

hartnn (hartnn):

sorry, x= 155/52

hartnn (hartnn):

u need help with question u posted or question u commented or both ?

OpenStudy (anonymous):

the question i posted i need to find x and y

OpenStudy (anonymous):

28x - 7 + 24 x + 18 = 180 52x + 11 = 180 52x = 169 x = 3.25 12y + 30 = 96 12 y = 66 x= 5.5 that's what i got and i think it's right, so thank you for helping me(:

hartnn (hartnn):

yup, correct.

OpenStudy (anonymous):

Thank you!

hartnn (hartnn):

u had one more ?